OpenStudy (anonymous):

SOME ONE PLEASE HELP ME! I'LL DO ANYTHING! Your portfolio must include a minimum of the following five types of equations and solutions: Two one-step equations Two equations that contains fractions One equation with distributive property One equation with decimals One real-world problem that is solved by an equation Once you have created each equation, you will solve it and show your work. Pretend that you are teaching the equations to a pre-algebra student.

OpenStudy (adajiamcneal):

this is easy -.- im pretty sure you know how to do this.

OpenStudy (here_to_help15):

lol

OpenStudy (anonymous):

like 1+1 is a one step eqaution

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

I could do it but I like have a ton of assignments left and I have to finish them all by the 21st this month..

OpenStudy (here_to_help15):

Connections Academy?

OpenStudy (anonymous):

FLVS

OpenStudy (anonymous):

i take FLVS what grade maybe i did the assignment

OpenStudy (anonymous):

8th

OpenStudy (anonymous):

lesson 7.06 right?

OpenStudy (anonymous):

yes @MadamNarwhal

OpenStudy (anonymous):

Which ones do u need cause some of them are really simple others i understand u may need help with

OpenStudy (anonymous):

atleast the fractions and decimals ones, i know the other ones

OpenStudy (anonymous):

okay give me a sec please

OpenStudy (anonymous):

Try to use diff # but u get the basic idea Example 1: r/3-2=1 r/3-2+2=1+2 r/3=3 r/3*3=3*3 r=9 (fraction

OpenStudy (anonymous):

is that one problem?

OpenStudy (anonymous):

nevrmind lol

OpenStudy (anonymous):

lol Again put diff # but here is the basic formula Example 7.5(3+x)=y 7.5(x)+7.5(3) 7.5x+22.5 7.5x-22.5+22.5-22.5 -15=x

OpenStudy (anonymous):

Im actually almost done the segment

OpenStudy (anonymous):

i know i should understand those prolbems, but honestly i dont...im so bad at math

OpenStudy (anonymous):

Lol thats how i feel about module 8. which one do u feel like you struggle with most

OpenStudy (anonymous):

everything, like it doesnt even make sense to me..i dont understand how the numbers come out to make a certain answer.

OpenStudy (anonymous):

Okay im HORRIBLE at explaining and i havent done this in a while...

OpenStudy (anonymous):

I have an idea, I'll make up the equations and you help me through them to get the answer.

OpenStudy (anonymous):

The only thing that i can think ( u will need this in module 7 and 8) of straight away is use the desmos graphing calculator and put your equation in and if both lines are exactly the same infinte answers if the two ines dont touch no aner and if the lies both touch at one point than one answer there will never be 2 answers thats one is just to throw u off

OpenStudy (anonymous):

again havent done this in probably a couple months

OpenStudy (anonymous):

Equations that have multiple steps need to be simplified before they can be solved. It's important to check for the distributive property, like terms, or moving variables. When it's necessary to move variables from one side of the equal sign to another, use addition or subtraction (the inverse operation) depending on the sign of the coefficient.

OpenStudy (anonymous):

Rational numbers are numbers that can be written as a fraction ab, where b is not 0. They can also be repeating decimals or terminating decimals. When rational numbers appear as fractions in equations, they can be removed by multiplying by the reciprocal or multiplicative inverse. A reciprocal is a fraction flipped upside down. It's important to simplify all other terms that need to be moved, using addition or subtraction prior to multiplying by the reciprocal. Problems with decimals can be solved using the strategy of inverse operations. When working with a fraction, do not attempt to simply convert it to a decimal. Solve the equation with the fraction in place. Sorry i know u have probaly already rad this but htis is probably all i can do tbh i have bad memory and unless im working on it than i really dont know what to do

OpenStudy (anonymous):

does this look good? @MadamNarwhal

OpenStudy (anonymous):

I know at least most of it are good examples im not sure about fractions (mostly becuase i haven done it i a while) but im confident in the rest,